Trains are connected by couplers, which are under tension as the locomotive pulls the train. Imagine you are on a train speeding up with a constant acceleration. As you move through the train from the locomotive to the last car, measuring the tension in each set of couplers, does the tension increase, decrease, or stay the same? When the engineer applies the brakes, the couplers are under compression. How does this compression force vary from the locomotive to the last car. (Assume only the brakes on the wheels of the engine are applied)
Answer #1

given acceleration of the train is constant a
now, let mass of each locomative be m, mass of coach be M and there be n locomotives
then
tension in the first coupler from the engine = T1
T1 = n*m*a
tension in second coupler = T2
T2 = (n-1)*m*a
hence for last coupler
an = m*a
hence
we can see the value of tension decreases as we move from the first locomotive to the last car

when brakes are applied, the decelertion is -a
hence
the compression in the couplers in this case decreases as wee move from the first locomotive to the last one